filament sensors sem to be stuck
My input 1 switch and the Homing Switch right before the splice core seem to be either stuck, or not sensing filament.
Input 1 will “see” the filament, but when I unload it, it still acts as if filament is loaded and is lit, and will not allow me to load new filament until I power cycle the Palette 2S.
the homing switch sensor will not detect any filament, causing it to continuously attempt to load filament even though filament is coming out the PTFE tube.
Is there a way to clean or troubleshoot these switches, or do I need to throw parts at it? I’ve already removed and reinstalled the switches to make sure it’s not a loose plug issue.
Is this a worthwhile discussion?